# Robotics meets Fluid Dynamics: A Characterization of the Induced Airflow below a Quadrotor as a Turbulent Jet

### v2

@@ I Introduction
Detailed computation of the induced flow around a drone is a challenging problem as it requires expensive and time-consuming computational fluid dynamics (CFD) simulations. To limit the computational demand, CFD simulations typically focus on millisecond time scales and simulate the airflow only close to the vehicle [17, 18, 19]. Such a simulation-based methodology is suitable to assess the drone flight performance, but not efficient in analyzing the far-field which extends over many meters away from the drone where one needs to rely on turbulence modeling [14, 16]. In addition, slow CFD simulations are not acceptable for a real-time planner or controller to dynamically coordinate a fleet of multiple quadrotors, such that the vehicles minimize interference with each other.
@@ II Related Work
For control and simulation tasks the primary goal of a model is to predict the motion of the vehicle given a certain actuation, that is, predict the forces and torques acting on its body. The most widespread models are rotor-based quadratic thrust and drag models [25, 26, 27] and more advanced models that are based on blade-element-momentum (BEM) theory which calculates the lift and drag at a propeller based on airfoil theory [28, 29, 30, 31, 32, 33, 34]. However, these approaches focus on estimating the aerodynamic forces and torques of the individual rotors but do not compute the combined total volumetric flow induced by the vehicle.
@@ IV-C Data Collection
We filter out all data points where the vehicle moves with a speed greater than 0.1 m/s to reduce the influence of transient effects. Despite being in a large indoor space, the anemometer registers a small ambient flow in the range of 0.06 m/s to 0.12 m/s. As this flow is already observed before the drone takes off, it is not primarily caused by recirculation effects but related to the testing facility. The background level is about 2 orders of magnitudes below the induced velocities and we correct for this static offset by subtracting the ambient airflow speed from the measurements.
After filtering we obtain around $10000\,\mathrm{s}$ of flight data (position and corresponding anemometer measurement) for each of the vehicles. For further processing, this scattered data is binned into gridded data as the median of all measurements within a grid cell.
@@ VI Results: Far Field
From the preliminary considerations of the airflow close to the vehicle, we see that the individual contributions from the quadrotor’s four propellers merge about 2.5-length scales below the drone. In this section, we present our novel experimental findings to answer the question:‘How much wind do the quadrotors make away from their propellers in the far field?’.
For the remainder of this section, cylindrical coordinates are used. The data is binned radially along $r$ and analyzed for downstream coordinates $s$, following the convention introduced in Fig. 2.
@@ VI-B Jet Flow Scaling
Next, we focus on the expansion of a turbulent jet which is described by eq. (6). For a turbulent jet, the flow expands linearly as a cone with approximately 12 deg opening angle [23]. Furthermore, the jet’s centerline velocity scales inversely proportional to the distance to the nozzle as described by eq. (5).
In Figure 8 (a) and (b) we find good agreement with these two key characteristics for the combined rotor flow of a quadrotor after $\tilde{s}=2.5$. Figure 8a and b show the normalized half-width $\tilde{r}_{1/2}$ and the normalized centerline velocity $\langle\tilde{U}_{\text{C}}\rangle$ of the radial velocity profiles as a function of the normalized distance $\tilde{s}$ to the drone’s rotor plane. The half-width and velocity have been obtained by fitting the velocity profile eq. (3) to the measurements at each $\tilde{s}$-slice for each drone. The triangles indicate datapoints located in the near-field (not considered for the fit) and circles indicate measurements from the far field (considered for the fit). The black lines represent a fit using the function form prescribed by the turbulent jet model eq. (5) and (6). The parameters obtained from the fit are
@@ VII-A Downwash Compensation
The electronic speed controllers (ESC) on the Kolibri drone exhibit a quadratic relationship between the throttle command and the power consumption. Consequently, the throttle setpoint computed by the controller is scaled with $\sqrt{\beta}$ before sending it to the vehicle’s motor controllers. This scaling factor is time-varying as it depends on $\Delta\mathbf{p}$. We implement eq. (11) on top of an MPC controller [46].
@@ VII-B Experimental Results
To demonstrate the effectiveness of the control scheme, we perform the following experiment: the Offboard 2 drones hovers at 3 m height and the much smaller Kolibri passes 1 and 2 meters exactly below at a translation speed of 1 length-scale per second. The results are shown in Fig. 10 and Table III and one can see that the baseline controller without the proposed downwash correction has a much higher altitude tracking error compared to the corrected controller. Comparing the RMSE in height we see a 3 to 5-fold improvement with the proposed controller and observe that the performance improvement is larger in stronger downwash (e.g., closer to the upper drones).

### v3

@@ Abstract
Video: https://youtu.be/-erfmxWTzPs
Index Terms:
Aerial Systems: Applications; Calibration and Identification; Robust/Adaptive Control
@@ I Introduction
Detailed computation of the induced flow around a drone is a challenging problem as it requires expensive and time-consuming computational fluid dynamics (CFD) simulations. To limit the computational demand, CFD simulations typically focus on millisecond time scales and simulate the airflow only close to the vehicle [17, 18, 19]. Such a simulation-based methodology is suitable to assess the drone flight performance, but not efficient in analyzing the far-field which extends over many meters away from the drone where one needs to rely on turbulence modeling [14, 16]. Additionally, slow CFD simulations are not usable in a real-time planner or controller to dynamically coordinate a fleet of quadrotors, such that the vehicles minimize interference with each other.
@@ II Related Work
For control and simulation tasks the primary goal of a model is to predict the motion of the vehicle given a certain actuation, that is, predict the forces and torques acting on its body. The most widespread models are rotor-based quadratic thrust and drag models [25, 26, 27] and more advanced models that are based on blade-element-momentum (BEM) theory which calculates the lift and drag at a propeller based on airfoil theory [28, 29, 30, 31, 32, 33, 34]. However, these approaches focus on estimating the aerodynamic forces and torques of the individual rotors but do not compute the combined total induced volumetric flow.
@@ IV-C Data Collection
We filter out all data points where the vehicle moves with a speed greater than 0.1 m/s to reduce the influence of transient effects. Despite being in a large indoor space, the anemometer registers a small ambient flow in the range of 0.06 m/s to 0.12 m/s. As this flow is already observed before the drone takes off, it is not primarily caused by recirculation effects but related to the testing facility. The background level is about 2 orders of magnitudes below the induced velocities and we correct for this static offset by subtracting the ambient airflow speed from the measurements. After filtering we obtain around $10000\,\mathrm{s}$ of flight data (position and corresponding anemometer measurement) for each of the vehicles. For further processing, this scattered data is binned into gridded data as the median of all measurements within a grid cell.
@@ VI Results: Far Field
From the preliminary considerations of the airflow close to the vehicle, we see that the individual contributions from the quadrotor’s four propellers merge about 2.5-length scales below the drone. In this section, we present our novel experimental findings to answer the question: ‘How much wind do the quadrotors make away from their propellers in the far field?’. For the remainder of this section, cylindrical coordinates are used. The data is binned radially along $r$ and analyzed for downstream coordinates $s$, following the convention introduced in Fig. 2.
@@ VI-B Jet Flow Scaling
Next, we focus on the expansion of a turbulent jet which is described by eq. (6). For a turbulent jet, the flow expands linearly as a cone with approximately 12 deg opening angle [23]. Furthermore, the jet’s centerline velocity scales inversely proportional to the distance as described by eq. (5).
In Figure 8 (a) and (b) we find good agreement with these two key characteristics for the combined rotor flow of a quadrotor after $\tilde{s}=2.5$. Figure 8a and b show the normalized half-width $\tilde{r}_{1/2}$ and the normalized centerline velocity $\langle\tilde{U}_{\text{C}}\rangle$ of the radial velocity profiles as a function of the normalized distance $\tilde{s}$ to the drone’s rotor plane. The half-width and velocity have been obtained by fitting the velocity profile eq. (3) to the measurements at each $\tilde{s}$-slice for each drone. The triangles indicate datapoints located in the near-field (not considered for the fit) and circles indicate measurements from the far field (considered for the fit). The black lines represent a fit using the function form prescribed by the turbulent jet model eq. (5) and (6). From the fit we obtain
@@ VII-A Downwash Compensation
The electronic speed controllers (ESC) on the Kolibri drone exhibit a quadratic relationship between the throttle command and the power consumption. Consequently, the throttle setpoint computed by the controller is scaled with $\sqrt{\beta}$ before sending it to the vehicle’s motor controllers. This scaling factor is time-varying as it depends on $\Delta\mathbf{p}$. We implement the downwash compensation on top of an MPC controller [46] by scaling the throttle command according to (11).
@@ VII-B Experimental Results
To demonstrate the effectiveness of the control scheme, we perform the following experiment: the Offboard 2 drones hovers at 3 m height and the much smaller Kolibri passes 1 and 2 meters exactly below at a translation speed of 1 length-scale per second. The results are shown in Fig. 10 and Table III and one can see that the baseline controller without the proposed downwash correction has a much higher altitude tracking error compared to the corrected controller. Directly below the drone the controller without correction deviates up to 30 cm from the reference height. Comparing the RMSE in height we see a 3 to 5-fold improvement with the proposed controller and observe that the performance improvement is larger in stronger downwash (e.g., closer to the upper drone).